A tennis ball rolls off the edge of a tabletop 0.75 π‘š above the floor and strikes the floor at a point 1.40 π‘š horizontally from the edge of the table. Ignore air resistance. a). Find the time of flight. b). Find the magnitude of the initial velocity. c). Find the magnitude and direction of the velocity of the ball just before it strikes the floor.


1
Expert's answer
2021-09-28T11:14:07-0400

(a) The time is


"t=\sqrt{\frac{2h}{g}}=0.39\text{ s}."

(b) The magnitude of the initial velocity is


"v_i=v_x=\frac Rt=1.9\text{ m/s}."

(c) The vertical component and the total velocity:


"v_y=\sqrt{2gh},\\
v=\sqrt{v_x^2+v_y^2}=\sqrt{v_x+2gh}=4.3\text{ m/s},\\\space\\
\theta=\arctan\frac{v_y}{v_x}=64°\text{ below the horizontal}."